if cos theta = 1/4, zero degrees < theta < 90degrees; what is csc theta??
Answer by lwsshak3(11628)

if cos theta = 1/4, zero degrees < theta < 90degrees; what is csc theta??
sin theta=√(1-cos^2 theta)=√(1-1/16)=√(15/16)=√15/4
csc theta=1/sin theta=4/√15=4√15/15